# Zirui ${}^{\clubsuit,\spadesuit}$   Zhizhou ${}^{\vardiamond,\spadesuit}$   Zheng ${}^{\spadesuit}$   Yilin ${}^{\vardiamond,\spadesuit}$   Zhuowen ${}^{\spadesuit}$

### v1

@@ Abstract
We present TokenCompose, a Latent Diffusion Model for text-to-image generation that achieves enhanced consistency between user-specified text prompts and model-generated images. Despite its tremendous success, the standard denoising process in the Latent Diffusion Model takes text prompts as conditions only, absent explicit constraint for the consistency between the text prompts and the image contents, leading to unsatisfactory results for composing multiple object categories. TokenCompose aims to improve multi-category instance composition by introducing the token-wise consistency terms between the image content and object segmentation maps in the finetuning stage. TokenCompose can be applied directly to the existing training pipeline of text-conditioned diffusion models without extra human labeling information. By finetuning Stable Diffusion, the model exhibits significant improvements in multi-category instance composition and enhanced photorealism for its generated images. [^1] Project done while Zirui Wang, Zhizhou Sha and Yilin Wang interned at UC San Diego. Project link: https://mlpc-ucsd.github.io/TokenCompose
@@ 1 Introduction
Despite the tremendous progress in recent text-to-image diffusion models [53, 56, 74, 58, 59] that have achieved creating images with an increasing level of quality, resolution, photorealism, and diversity, there still exists a major consistency problem between the text prompt and the generated image content. The models loose the composition capability when multiple object categories, especially those not commonly appearing simultaneously in the real world, are included in the text prompt: objects may not appear in the image or their configuration is not pleasantly good looking. Figure 1 shows examples where a state-of-the-art model, Stable Diffusion [56], fails to generate desirable image content from text prompts.
@@ 2 Related Works
Generative Models for Image Understanding. There has been an upward trend in the use of text-conditioned image generative models for open-vocabulary image understanding tasks, such as classification [34, 11], detection [36, 8], and segmentation [38, 70, 65, 29, 45, 49, 68, 69, 64].
@@ 3.1 Preliminaries
$\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\Big{]}\,$ (1)
@@
$\mathcal{L}_{LDM}:=\mathbb{E}_{\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))\|_{2}^{2}\Big{]}\,$ (2)
@@ 3.2 Token-level Attention Loss
$Q^{(h)}=W^{(h)}_{Q}\cdot\varphi(z_{t}),\;K^{(h)}=W^{(h)}_{K}\cdot\tau_{\theta}(y)\;$ (3)
$\mathcal{A}=\frac{1}{H}\sum_{h}^{H}\text{softmax}\left(\frac{Q^{(h)}(K^{(h)})^{T}}{\sqrt{d_{k}}}\right)$ (4)
@@
Figure 3: Illustration of $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$. We illustrate how $\mathcal{L}_{\text{token}}$ (bottom-right) and $\mathcal{L}_{\text{pixel}}$ (upper-right) are calculated given a cross-attention map $\mathcal{A}_{i}$ and a binary segmentation mask $\mathcal{M}_{i}$. $\mathcal{L}_{\text{token}}$ aggregates attention activations toward non-masked regions of $\mathcal{M}_{i}$, and this objective is normalized by the total activations of $\mathcal{A}_{i}$. However, it does not constrain where activations should be once inside the non-masked region. $\mathcal{L}_{\text{pixel}}$ gives precise supervision whether a pixel belongs to the segmented region, constraining where activations should be with binary values. However, it is not normalized by the total activations of $\mathcal{A}_{i}$. Combining $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, we take advantage of the benefit of each objective while minimizing their side effects to a minimum level. We show examples of cross-attention activations from models optimized with $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, either of them, and neither of them in Figure 4.
@@ 3.3 Pixel-level Attention Loss
Figure 4: Impact on cross-attention activations with different objectives. We firstly demonstrate that finetuning the Stable Diffusion with only $\mathcal{L}_{LDM}$ does not improve grounding capabilities as much. Adding $\mathcal{L_{\text{pixel}}}$ alone causes increased cross-attention activations in general. Adding $\mathcal{L_{\text{token}}}$ plays a vital role in improving token grounding, but leads activations to aggregate in subregions of the targets. By combining $\mathcal{L_{\text{token}}}$ and $\mathcal{L_{\text{pixel}}}$, the model shows substantial improvement in grounding text tokens with image features. In this illustration, we apply the null text inversion [48] technique to all models, allowing them to generate the same image for comparable cross-attention maps. In addition, we decompose cross-attention map visualizations by different layers, attention heads, and denoising timesteps, and the results are shown in Appendix E.
@@ 4 Experiment
Multi-category Instance Composition ()normal-↑($\uparrow$)( ↑ ) Photorealism ()normal-↓($\downarrow$)( ↓ ) Eff. ()normal-↓($\downarrow$)( ↓ )
@@ 4.1 Training Details
Dataset. To study the effectiveness of token grounding objectives, we finetune the Stable Diffusion model on a subset of COCO image-caption pairs [39]. Specifically, we first select all unique images from the Visual Spatial Reasoning [40] dataset, as these images have fewer visual-linguistic ambiguities and a greater number of different categories that appear in each image. Then, we use a CLIP model [52] to select the caption with the highest semantic similarity with respect to its corresponding image. Finally, we adopt a pretrained noun parser to parse all nouns from the captions, and leverage Grounded-SAM to generate binary segmentation masks for each noun (or noun phrase) [1, 31, 42]. The final dataset consists of $\sim$ 4526 image-caption pairs and their respective binary segmentation masks. We illustrate a high-level data and training pipeline in Figure 2. We provide visualizations of dataset samples in Appendix C.
Setup. Our main experiments are performed using Stable Diffusion v1.4 [56], a popular text-to-image diffusion model for high-quality generation. We use both $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ in addition to its original denoising objective $\mathcal{L}_{LDM}$ with a constant global learning rate of 5e-6 for 24,000 steps (32,000 for Stable Diffusion v2.1) using the AdamW optimizer [44]. We trained the entire U-Net with a batch size of 1 and 4 gradient accumulation steps on a single GPU. We apply center crop to all training images and their respective segmentation maps $\mathcal{M}$. For Stable Diffusion v1.4, the cross-attention layers are located in the U-Net encoder $U_{\text{E}}$, the middle block $U_{\text{Mid}}$, and the decoder $U_{\text{D}}$. We apply $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ to all cross-attention layers in $U_{\text{Mid}}$ and $U_{\text{D}}$, but not $U_{\text{E}}$. We provide detailed setup for the entire pipeline in Appendix A.
@@ 4.2 Main Results
Multi-category Instance Composition. Benchmarks that examine this compositionality focus on the successful generation of multiple categories of instances in the image mentioned in the text condition. We use an existing benchmark, VISOR [17], along with our benchmark, MultiGen, to study the model’s capability in multi-category instance composition. The VISOR benchmark obtains all unique pairwise combinations of the 80 object categories from COCO [39] and converts each pair (A, B) into a text prompt with an arbitrary spatial relationship (R) following a template “$<$A$>$ $<$R$>$ $<$B$>$”, for example, “a motorcycle to the left of an elephant.” With the images generated from such prompts, VISOR uses an open-vocabulary detector [46] to detect the presence and spatial locations of each category in the pair. Object Accuracy (OA) measures the successful rate of generating instances from both categories. VISOR also provides metrics for spatial relationships. However, since our work focuses on multi-category instance composition, we only adopt the OA metric, and report relevant numbers in Table 1.
@@
Compositional image generation often involves inference variance. To account for this, each prompt in MultiGen is used to generate 10 rounds of images, which results in 10 $\times$ 1000 images being generated for each dataset’s category combinations. We calculate MG2 through 5 for each round and report the mean and standard deviation (in subscript) of the MG2-5 success rate out of 10 rounds in Table 1. Based on the evaluation, our model exceeds all baselines in object accuracy and MG2-4. We find that Attend-and-Excite [6] has a considerable success rate in generating all 5 categories, but it falls short of generating 2-4 categories compared to our model. We conjecture that this is due to the training distribution where captions do not include as many as 5 or more categories of instances, which inevitably leads to diminishing improvements in multi-category instance composition. In addition to aggregated results, we provide further analysis of success rate of difference categories of instances in Appendix F.
Photorealism. We compare the image quality generated from the baselines and our model using the Fréchet Inception Distance (FID) metric [21]. For baselines that do not require additional conditional input other than captions, we calculate the FID score based on 10,000 image-caption pairs sampled from the COCO validation set (C). We also report the FID metric based on 1,000 image-caption pairs from the validation set of Flickr30K entities (F) [51]. As this dataset provides labels and bounding boxes for entities in the captions [72], it can be used by inference-based methods that require such input. We report all applicable scores in Table 1. We also qualitatively evaluate image quality in conjunction with multi-category instance composition across different baselines compared to our model in Figure 5 and in Appendix D. For fairness, we use the same initial latent in each comparison.
Efficiency. As a training-based method, our model does not require additional inference-time manipulations compared to a standard generation pipeline of a text-conditioned diffusion model. On the other hand, the majority of inference-based methods impose a non-ignorable compute burden for compositional generation, where the slowest baseline, Attend-and-Excite [6], takes more than 3$\times$ of time to generate a single image. We report efficiency results in Table 1 in seconds needed to generate an image on a single NVIDIA RTX 3090 GPU with 50 DDIM steps [63] and classifier-free guidance [23].
@@ 4.3 Generalization
Multi-category Instance Composition ()normal-↑($\uparrow$)( ↑ )
@@ 4.5 Downstream Metrics
We believe that improvements in these metrics are due to the model having a higher chance of composing multiple categories of instances successfully, which serves as a prerequisite for assigning attributes to and relations between objects. We provide further analysis in Appendix B.
Attribute Binding ($\uparrow$) Object Relations (normal-$\uparrow$)
@@ 5.2 Layers with Grounding Objectives
Finally, we experiment with adding $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$ at different layers of cross-attention of the denoising U-Net (v). We find that adding grounding objectives to the middle block and the decoder of the U-Net improves the overall performance of multi-category instance composition. Removing the constraint from the middle block or adding the constraint to the encoder degrades the performance of the model. Furthermore, for cross-attention layers at the decoder with variable resolutions, we find that the more layers optimized with $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$, the better the performance in both multi-category instance composition and photorealism.
@@ Acknowledgement
This work is supported by NSF Award IIS-2127544. We thank Yifan Xu, Dou Kwark, and Kaiyi Huang for their valuable feedback.
@@ G. DINO Model groundingdino_swint_ogc [42, 5, 43]
B Conditional Downstream Metrics
@@
C Grounded COCO Dataset
As seen in Figure 2, we adopt a POS Tagger and the Grounded SAM [1, 42, 31] to extract binary segmentation maps from noun tokens from the image-text pair dataset. In this section, we aim to expand the visualizations of the generated data in Figure 8.
Each row and column in Figure 8 represents a single data that the model is trained with. From the left to the right of the data are the caption, the input image, and the grounded binary segmentation maps.
The bold and underlined text in the caption represents noun tokens captured by the POS tagger where their corresponding segmentation maps are extracted from the Grounded SAM. Italicized and red text in the caption represents noun tokens captured by the POS tagger but does not have segmentation maps extracted from the Grounded SAM due to the model not being able to find the corresponding objects.
@@ C Grounded COCO Dataset
D Sampled Images from MultiGen
MultiGen benchmark plays a crucial role in evaluating multi-category instance composition capabilities of text-to-image models. Our model exhibits significant improvements in multi-category instance composition compared to previous methods, as evidenced by the quantitative results from Table 1.
To offer a more intuitive sense in this compositional task, we visualize the generated images from various prompts in this benchmark (e.g., MG COCO Instances and MG ADE20K Instances). We use the same latent for each comparison for fairness. The images are presented in Figure 9.
E Attention Map Visualizations

### v2

# https://mlpc-ucsd.github.io/TokenCompose

@@ Abstract
We present TokenCompose, a Latent Diffusion Model for text-to-image generation that achieves enhanced consistency between user-specified text prompts and model-generated images. Despite its tremendous success, the standard denoising process in the Latent Diffusion Model takes text prompts as conditions only, absent explicit constraint for the consistency between the text prompts and the image contents, leading to unsatisfactory results for composing multiple object categories. Our proposed TokenCompose aims to improve multi-category instance composition by introducing the token-wise consistency terms between the image content and object segmentation maps in the finetuning stage. TokenCompose can be applied directly to the existing training pipeline of text-conditioned diffusion models without extra human labeling information. By finetuning Stable Diffusion with our approach, the model exhibits significant improvements in multi-category instance composition and enhanced photorealism for its generated images.  [^1] Project done while ZW, ZS and YW interned at UC, San Diego. Correspondence to ZW at zw1300@cs.princeton.edu
@@ 1 Introduction
Despite the tremendous progress in recent text-to-image diffusion models [56, 59, 79, 61, 62, 18, 76, 7, 22] that have achieved creating images with an increasing level of quality, resolution, photorealism, and diversity, there still exists a major consistency problem between the text prompt and the generated image content. The models loose the composition capability when multiple object categories, especially those not commonly appearing simultaneously in the real world, are included in the text prompt: objects may not appear in the image or their configuration is not pleasantly good looking. Figure 1 shows examples where a state-of-the-art model, Stable Diffusion [59], fails to generate desirable image content from text prompts.
@@ 2 Related Works
Generative Models for Image Understanding. There has been an upward trend in the use of text-conditioned image generative models for open-vocabulary image understanding tasks, such as classification [37, 12], detection [39, 9], and segmentation [41, 74, 68, 32, 48, 52, 71, 73, 67, 72, 81].
@@ 3.1 Preliminaries
$\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\Big{]}\,\vspace{-1mm}$ (1)
@@
$\mathcal{L}_{LDM}:=\mathbb{E}_{\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))\|_{2}^{2}\Big{]}\,\vspace{-2mm}$ (2)
@@ 3.2 Token-level Attention Loss
$Q^{(h)}=W^{(h)}_{Q}\cdot\varphi(z_{t}),\;K^{(h)}=W^{(h)}_{K}\cdot\tau_{\theta}(y)\;\vspace{-2mm}$ (3)
$\mathcal{A}=\frac{1}{H}\sum_{h}^{H}\text{softmax}\left(\frac{Q^{(h)}(K^{(h)})^{T}}{\sqrt{d_{k}}}\right)\vspace{-2mm}$ (4)
@@
Figure 3: Illustration of $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$. We illustrate how $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ are calculated given a cross-attention map $\mathcal{A}_{i}$ and a binary segmentation mask $\mathcal{M}_{i}$. $\mathcal{L}_{\text{token}}$ aggregates attention activations toward non-masked regions of $\mathcal{M}_{i}$, and this objective is normalized by the total activations of $\mathcal{A}_{i}$. However, it does not constrain where activations should be once inside the non-masked region. $\mathcal{L}_{\text{pixel}}$ gives precise supervision whether a pixel belongs to the segmented region, constraining where activations should be with binary values. However, it is not normalized by the total activations of $\mathcal{A}_{i}$. Combining $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, we take advantage of the benefit of each objective while minimizing their side effects to a minimum level. We show examples of cross-attention activations from models optimized with $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, either of them, and neither of them in Figure 4.
@@ 3.3 Pixel-level Attention Loss
Figure 4: Impact on cross-attention activations with different objectives. We firstly demonstrate that finetuning the Stable Diffusion with only $\mathcal{L}_{LDM}$ does not improve grounding capabilities as much. Adding $\mathcal{L_{\text{pixel}}}$ alone causes increased cross-attention activations in general. Adding $\mathcal{L_{\text{token}}}$ plays a vital role in improving token grounding, but leads activations to aggregate in subregions of the targets. By combining $\mathcal{L_{\text{token}}}$ and $\mathcal{L_{\text{pixel}}}$, the model shows substantial improvement in grounding text tokens with image features. In this illustration, we apply the null text inversion [51] technique to all models, allowing them to generate the same image for comparable cross-attention maps.
@@ 4 Experiment
Multi-category Instance Composition ()($\uparrow$)( ↑ ) Photorealism ()($\downarrow$)( ↓ ) Eff. ()($\downarrow$)( ↓ )
@@ 4.1 Training Details
Dataset. To study the effectiveness of token grounding objectives, we finetune the Stable Diffusion model on a subset of COCO image-caption pairs [42]. Specifically, we first select all unique images from the Visual Spatial Reasoning [43] dataset, as these images have fewer visual-linguistic ambiguities and a greater number of different categories that appear in each image. Then, we use a CLIP model [55] to select the caption with the highest semantic similarity with respect to its corresponding image. Finally, we adopt a pretrained noun parser to parse all nouns from the captions, and leverage Grounded-SAM to generate binary segmentation masks for each noun (or noun phrase) [1, 34, 45]. The final dataset consists of $\sim$ 4526 image-caption pairs and their respective binary segmentation masks. We illustrate a high-level data and training pipeline in Figure 2.
Setup. Our main experiments are performed using Stable Diffusion v1.4 [59], a popular text-to-image diffusion model for high-quality generation. We use both $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ in addition to its original denoising objective $\mathcal{L}_{LDM}$ with a constant global learning rate of 5e-6 for 24,000 steps (32,000 for Stable Diffusion v2.1) using the AdamW optimizer [47]. We trained the entire U-Net with a batch size of 1 and 4 gradient accumulation steps on a single GPU. We apply center crop to all training images and their respective segmentation maps $\mathcal{M}$. For Stable Diffusion v1.4, the cross-attention layers are located in the U-Net encoder $U_{\text{E}}$, the middle block $U_{\text{Mid}}$, and the decoder $U_{\text{D}}$. We apply $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ to all cross-attention layers in $U_{\text{Mid}}$ and $U_{\text{D}}$.
@@ 4.2 Main Results
Multi-category Instance Composition. Benchmarks that examine this compositionality focus on the successful generation of multiple categories of instances in the image mentioned in the text condition. We use an existing benchmark, VISOR [19], along with our benchmark, MultiGen, to study the model’s capability in multi-category instance composition. The VISOR benchmark obtains all unique pairwise combinations of the 80 object categories from COCO [42] and converts each pair (A, B) into a text prompt with an arbitrary spatial relationship (R) following a template “$<$A$>$ $<$R$>$ $<$B$>$", for example, “a motorcycle to the left of an elephant." With the images generated from such prompts, VISOR uses an open-vocabulary detector [49] to detect the presence and spatial locations of each category in the pair. Object Accuracy (OA) measures the successful rate of generating instances from both categories. VISOR also provides metrics for spatial relationships. However, since our work focuses on multi-category instance composition, we only adopt the OA metric, and report relevant numbers in Table 1.
@@
Compositional image generation often involves inference variance. To account for this, each prompt in MultiGen is used to generate 10 rounds of images, which results in 10 $\times$ 1000 images being generated for each dataset’s category combinations. We calculate MG2-5 for each round and report the mean and standard deviation (in subscript) of the MG2-5 success rate out of 10 rounds in Table 1. Based on the evaluation, our model exceeds all baselines in object accuracy and MG2-4. We find that Attend-and-Excite [6] has a considerable success rate in generating all 5 categories, but it falls short of generating 2-4 categories comparably. We conjecture that this is due to the training distribution where captions do not include as many as 5 or more categories of instances, which inevitably leads to diminishing improvements in multi-category instance composition.
Photorealism. We compare the image quality generated from the baselines and our model using the Fréchet Inception Distance (FID) metric [24]. For baselines that do not require additional conditional input other than captions, we calculate the FID score based on 10,000 image-caption pairs sampled from the COCO validation set (C). We also report the FID metric based on 1,000 image-caption pairs from the validation set of Flickr30K entities (F) [54]. As this dataset provides labels and bounding boxes for entities in the captions [77], it can be used by inference-based methods that require such input. We report all applicable scores in Table 1. We also qualitatively evaluate image quality in conjunction with multi-category instance composition across different baselines compared to our model in Figure 5. For fairness, we use the same initial latent in each comparison.
Efficiency. As a training-based method, our model does not require additional inference-time manipulations compared to a standard text-to-image diffusion pipeline. On the other hand, the majority of inference-based methods impose a non-ignorable compute burden for compositional generation, where the slowest baseline, Attend-and-Excite [6], takes more than 3$\times$ of time to generate a single image. We report efficiency results in Table 1 in seconds needed to generate an image on a single NVIDIA RTX 3090 GPU with 50 DDIM steps [66] and classifier-free guidance [26].
@@ 4.3 Generalization
Multi-category Instance Composition ()($\uparrow$)( ↑ )
@@ 4.5 Downstream Metrics
We believe that improvements in these metrics are due to the model having higher chances of composing multiple categories of instances, which serves as a prerequisite for assigning attributes to and relations between objects.
Attribute Binding ($\uparrow$) Object Relations ($\uparrow$)
@@ 5.2 Layers with Grounding Objectives
Finally, we experiment with adding $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$ at different layers of cross-attention of the denoising U-Net (v). We find that adding grounding objectives to the middle block and the decoder of the U-Net improves the overall performance of multi-category instance composition. Removing the constraint from the middle block or adding the constraint to the encoder degrades the performance. Furthermore, for cross-attention layers at the decoder with variable resolutions, we find that the more layers optimized with $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$, the better the performance in both multi-category instance composition and photorealism.
@@ Acknowledgement
This work is supported by NSF Award IIS-2127544. We thank Yifan Xu and Dou Kwark from UC San Diego, Kaiyi Huang from University of Hong Kong, and Adithya Bhaskar and Ofir Press from the Princeton NLP Group for discussions and/or feedback.
@@ G. DINO Model groundingdino_swint_ogc [45, 5, 46]
8 Conditional Downstream Metrics
@@
Figure 8: More samples in multi-category instance composition.
Figure 9: Downstream applications in prompt-to-prompt [22] image editing and zero-shot outpainting and inpainting.
Figure 10: Categories with poor & strong performance
Figure 11: Failure case analysis in multi-category compositionality.
9 More Examples
Multi-category Instance Composition. We provide more visualizations in Figure. 8 to illustrate capabilities of TokenCompose in multi-category instance composition. In addition to being able to generate multiple categories of instances successfully, we show that object affordance (e.g., attach, sit, support, etc) can be maintained, which indicates that TokenCompose is able to implicitly learn the basic “physical rules" via the object token and segmentation consistency constraints.
Downstream Applications. We show in Figure. 9 positive impacts in downstream applications. As expected, TokenCompose shows its effectiveness in tasks beyond text-to-image generation.
Benchmark. MultiGen benchmark plays a crucial role in evaluating multi-category instance composition capabilities of text-to-image models. To offer a more intuitive sense in this compositional task, we visualize the generated images from various prompts in this benchmark (e.g., MG COCO Instances and MG ADE20K Instances). We use the same latent for each comparison for fairness. The images are presented in Figure 12.
10 Grounded COCO Dataset
As seen in Figure 2, we adopt a POS Tagger and the Grounded SAM [1, 45, 34] to extract binary segmentation maps from noun tokens from the image-text pair dataset. We aim to expand the visualizations of the generated data in Figure 13. Each row and column represents a single data that the model is trained with. From the left to the right of the data are the caption, the input image, and the grounded binary segmentation maps.
The bold and underlined text in the caption represents noun tokens captured by the POS tagger where their corresponding segmentation maps are extracted from the Grounded SAM. Italicized and red text in the caption represents noun tokens captured by the POS tagger but does not have segmentation maps extracted from the Grounded SAM due to the model not being able to locate the objects.
@@ 10 Grounded COCO Dataset
11 Analysis
Attention Visualizations. To gain a better understanding of how incorporating grounding objectives into text-to-image models during training affects cross-attention maps for image reconstruction [51] & generation tasks at inference time, we provide token-level cross-attention map visualizations on three axes: (1) different cross-attention layers with various resolutions (Figure 14); (2) different heads of the multi-head cross-attention (Figure 15; and (3) different timesteps during the denoising process (Figure 16).
Multi-category Instance Composition Success Rate. Given a set of compositional prompts, we calculate the count of each category that appears in these prompts along with images where the instance(s) of this category is/are detected by a detection model [50]. We divide the number of images that contain a specific category of the instance by the number of prompts that contains this category to acquire the success rate and report the numbers in Figure 17.
Failure Cases. We provide generated and detected samples for categories with low and high success rates in Figure. 10. We believe that one explanation for the poor performance categories have variants and viewpoints with drastic visual differences, so learning and generating them is harder. Further, we aggregate patterns of common failures in our multi-category instance composition in Figure. 11. We believe that visual commonsense reasoning aspect of the generative model would be an area of improvement.